Complete the truth tables for these two Boolean expressions: (a) Output = A + B (b) Output = A + AB


1
Expert's answer
2021-11-22T16:20:28-0500

Let us complete the truth tables for these two Boolean expressions:


(a) Output = A + B


"\begin{array}{||c|c||c||}
\hline\hline
A & B & A+B\\
\hline\hline
0 & 0 & 0\\
\hline
0 & 1 & 1 \\
\hline
1 & 0 & 1\\
\hline
1 & 1 & 1\\
\hline\hline
\end{array}"


(b) Output = A + AB


"\begin{array}{||c|c||c|c||}
\hline\hline
A & B & AB& A+AB\\
\hline\hline
0 & 0 & 0 & 0\\
\hline
0 & 1 & 0 & 0\\
\hline
1 & 0 & 0 & 1\\
\hline
1 & 1 & 1 & 1\\
\hline\hline
\end{array}"